OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 -7 7 9 -9

OpenStudy (anonymous):

Multiply first equation with 4 and second with 5: 20x + 16y = 4 20x + 15 y = -5 Subtract them: y = 4-(-5) = 4 + 5 = ?? Can you do it now? Just to add 4 and 5..

OpenStudy (anonymous):

thank you! just so i did it right can you show me your answer so i can see what i did wrong

OpenStudy (anonymous):

Firstly tell me yours..

OpenStudy (anonymous):

y=29

OpenStudy (anonymous):

What will you get when you will add 4 and 5???

OpenStudy (anonymous):

.....ohh sorry 9

OpenStudy (anonymous):

Yes so this is the answer what is problem in that??

OpenStudy (anonymous):

..what?